# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ Abstract.
Our main contributions are threefold. First, we establish a decomposition theorem, proving that any such projection-valued map admits a splitting into two projection-valued maps that are related to each other via time-reversal symmetry. Second, we provide a complete homotopy classification theorem for these maps, thereby clarifying their topological structure. Third, by means of the previous analysis, we connect the Fu–Kane–Mele index to the Chern number of one of the factors in the previously-mentioned decomposition, which in turn allows to exhibit how the $\mathbb{Z}_{2}$-valued topological obstruction to constructing a periodic and smooth Bloch frame for the projection-valued map, measured by the Fu–Kane–Mele index, can be concentrated in a single pseudo-periodic Kramers pair.
@@ Introduction
Class AII corresponds to systems endowed with an antiunitary time-reversal symmetry $\Theta$ satisfying $\Theta^{2}=-1$. For crystalline and insulating systems, described by a periodic Hamiltonian with a spectral gap, one associates the so-called Fermi projector $P$, which defines the occupied subspace. This, in turn, gives rise to the Bloch bundle $E\to\mathbb{T}^{d}$ over the Brillouin torus $\mathbb{T}^{d}$, carrying an induced antiunitary equivariant structure. Fiber bundles subject to this type of symmetry are called “Quaternionic”-vector bundles, and have became object of independent geometric interest, beyond the specific case considered in this paper, due to their peculiar topological structure (see, e.g., [de2015classification, bunk2019topologicalinsulatorskanemeleinvariant]).
As shown in [panati2007triviality, monaco2015symmetry], the presence of $\Theta$ guarantees the existence of a global smooth trivialization of $E$. However, such a trivialization need not respect the time-reversal symmetry itself, and the obstruction to the existence of a symmetric trivialization carries non-trivial topological content.
@@
In two-dimensional systems, this obstruction is quantified by the Fu–Kane–Mele (FKM) invariant, first introduced in [kane2005z]. While the original formulation of this invariant was already deeply topological in nature, several alternative formulations and interpretations have since been proposed, both in physical [FuKane2006, MooreBalents2007] and in mathematical terms, see e.g. [graf2013bulk, de2015classification, fiorenza2016z, monaco2015symmetry]. For an overview of the relation among some of the different approaches, we refer the reader to [Cornean_2017].
@@
Following the approach initiated in [fiorenza2016z], the $\mathbb{Z}_{2}$ invariant can be interpreted as the obstruction to the existence of a globally symmetric trivialization of the Bloch bundle. In position space, the analog of a globally smooth trivialization of the Bloch bundle is the existence of exponentially localized Wannier bases. Such bases have a clear physical interpretation as representing localized electronic orbitals, and their existence is intimately related to the insulating nature of the system. The requirement that a Wannier basis be symmetric under time-reversal symmetry allows to implement the symmetry at the level of such orbitals, which can be useful when these are used, for example, to derive symmetric tight-binding models from continuum Hamiltonians [marzari2012maximally].
@@
If, in addition to $\Theta$, the system possesses a unitary involutive symmetry $S$ commuting suitably with $\Theta$—corresponding, for example, to spin conservation—then the topological invariant persists and admits a simple physical interpretation in position space as the spin Chern number [Sheng_2006, Prodan_2009]. In this case, $S$ splits the Fermi projector into two components,
(1) $P_{\pm}:=\frac{1}{2}\big(P\pm PS\big),$
which are exchanged by $\Theta$, and the topological invariant corresponds to the Chern number of either component. As shown by Prodan [Prodan_2009], the spin Chern number is stable under small perturbations, remaining well-defined modulo $2$ even when the symmetry $S$ is slightly broken—thus capturing the $\mathbb{Z}_{2}$ topology of class AII systems.
@@
In the present work, we prove that, for periodic systems endowed with fermionic time-reversal symmetry, it is always possible to decompose the Fermi projector $P$ into a pair of projectors $P_{+}$ and $P_{-}$, related to each other under time reversal symmetry, in direct analogy with the spin-conserving case. We show that the topological characterization of such systems is captured by the Chern number, modulo $2$, of one component of the decomposition. The $\mathbb{Z}_{2}$ nature of the invariant arises from the non-uniqueness of this decomposition. This result provides a new perspective on the geometric and physical meaning of the Fu–Kane–Mele index: unlike other topological formulations, it admits a clear position-space counterpart, namely the Chern marker of one of the decomposition projectors. This position-space counterpart of the invariant paves the way for extending several results previously obtained in the spin-conserving case to the more general symmetry class AII, such as the Středa formula [Monaco_2020].
The paper is structured as follows. The first section is devoted to describing the standard procedure by which, starting from a symmetric and periodic Hamiltonian, one obtains the object at the center of our mathematical study: periodic projection-valued maps. The second section revisits parallel transport and the first Chern number in the context of continuous, rather than differentiable, projection-valued maps. Although this section does not contain original results, we believe it is useful to provide a concise and coherent presentation of these techniques.
The third section introduces our definition of the $\mathbb{Z}_{2}$-valued topological invariant, which differs from previous definitions but, at first glance, offers no computational or interpretative advantage. The fourth section marks our first original and physically relevant contribution, as it establishes an equivalence between the previously defined invariant and the possibility of decomposing the considered projection-valued map into a pair of symmetric maps whose Chern numbers have parity consistent with the invariant.
In the fifth section, we study the problem of constructing sections for a projection-valued map. In particular, this problem was previously approached by interpreting the $\mathbb{Z}_{2}$ invariant as an obstruction to realizing a periodic, continuous, and symmetric section. Our approach, instead, interprets the invariant as a set of gluing conditions for a section that need not be periodic.
The sixth and final section confirms the completeness of the invariant with respect to unitary equivalence and homotopy relations. The article concludes with an appendix presenting a series of classical results in algebraic topology that are used throughout the paper.
1. Setting of the problem: from periodic Hamiltonians to projection-valued maps
@@
In this section, we outline the standard procedure that allows one to associate a projection-valued map to a periodic Hamiltonian, following the construction reviewed in [panati2007triviality, panati2013bloch] to which we refer the reader for further details. We consider a quantum particle moving in a condensed matter system, whose configuration space $X\subset\mathbb{R}^{d}$ exhibits a periodic structure with respect to a lattice $\Lambda\simeq\mathbb{Z}^{d}$, commonly referred to as the Bravais lattice. Explicitly, this periodicity implies $X=\lambda+X$ for every $\lambda\in\Lambda$. Two types of models are typically used in this context: continuum models, in which $X=\mathbb{R}^{d}$, and tight-binding models, where $X$ is replaced by the discrete set of atomic positions and the quotient $X/\Lambda$ is finite. When additional internal degrees of freedom, such as spin, are present, the configuration space is naturally extended to $X\times\{1,\ldots,n\}$, and the corresponding Hamiltonian $H$ describing the system acts on $L^{2}(X)\otimes\mathbb{C}^{n}$. Furthermore, the system is assumed to be endowed with a fermionic time-reversal symmetry, i.e., an anti-unitary operator $\Theta$ acting on $L^{2}(X)\otimes\mathbb{C}^{n}$ such that $\Theta^{2}=-\mathds{1}$, commuting with the Hamiltonian $H$.
For the sake of presentation, in the following we focus on continuum models ($X=\mathbb{R}^{d}$), having in mind time-reversal symmetric systems modelled by Hamiltonians with spin-orbit coupling of Rashba type [rashba2005spin], but all that will be discussed applies just as well to tight-binding models, like the paradigmatic Kane–Mele Hamiltonian [Kane_2005, kane2005z].
@@
where $(T_{\lambda}\phi)(x)=\phi(x-\lambda)$ for all $\lambda\in\Lambda$. Correspondingly, the fermionic time-reversal symmetry operator $\Theta$ is required to be itself translation-invariant, i.e., to commute with the lattice translations $T_{\lambda}$.
A powerful tool to study the spectral properties of such periodic Hamiltonians is the Bloch-Floquet-Zak transform [kuchment2016overview], which allows one to decompose the Hilbert space into fibers labeled by the quasi-momentum $k$. For any sufficiently rapidly decaying $\phi\in L^{2}(\mathbb{R}^{d})$, the transform is defined as the following function depending on $k\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$:
@@
$(\mathcal{U}_{\mathrm{BFZ}}\phi)_{k}(y)\in\left\{f\in L^{2}_{\rm loc}(\mathbb{R}^{d},dk,\mathcal{H}_{f})\ \middle|\ f(k+\lambda^{*})=\tau_{\lambda^{*}}f(k)\right\}\simeq\int_{k\in\mathbb{B}}^{\oplus}dk\,\mathcal{H}_{f},$
where $\mathcal{H}_{f}$ is the space of $\Lambda$-periodic, locally square-integrable functions on $\mathbb{R}^{d}$, and $\tau_{\lambda^{*}}:\mathcal{H}_{f}\to\mathcal{H}_{f}$ acts as $\tau_{\lambda^{*}}(g)(y)=e^{i\lambda^{*}\cdot y}g(y)$. By also choosing a periodicity cell for $\Lambda$, the Wigner-Seitz cell $\mathbb{W}$, one can further identify $\mathcal{H}_{f}$ with $L^{2}(\mathbb{W})$. A classical result, which can be traced back to [bloch1929quantenmechanik, zak1964magnetic], reads:
@@
The operator $\mathcal{U}_{\mathrm{BFZ}}$ extends uniquely to a unitary operator
$\mathcal{U}_{\mathrm{BFZ}}:L^{2}(\mathbb{R}^{d})\longrightarrow\int_{k\in\mathbb{B}}^{\oplus}dk\,L^{2}_{\rm per}(\mathbb{W}).$
When $H$ acts on $L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{n}$, the transformed operator $(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})H(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}$, acting on
$\mathcal{H}=\left(\int_{k\in\mathbb{B}}^{\oplus}dk\,L^{2}_{\rm per}(\mathbb{W})\right)\otimes\mathbb{C}^{n}\simeq\int_{k\in\mathbb{B}}^{\oplus}dk\,\left(L^{2}_{\rm per}(\mathbb{W})\otimes\mathbb{C}^{n}\right),$
is decomposable. More precisely, there exists a family of self-adjoint fiber operators $H(k)$ such that
$\big[(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})H(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}f\big]_{k}=H(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{R}^{d},$
@@
The same procedure can be applied to $\Theta$, since it commutes with the translation operators. Consequently, there exists a family of anti-unitary operators $T(k)$ acting on $\mathcal{H}$ such that
$\left[(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})\Theta(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}(f)\right]_{k}=T(k)f_{-k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{R}^{d}.$
@@
$T(k+\lambda^{*})=\tau_{\lambda^{*}}T(k)\tau_{\lambda^{*}},\quad\forall k\in\mathbb{B},\ \lambda^{*}\in\Lambda^{*}.$
In most physical models, $\Theta$ acts only on the second factor of the tensor product $L^{2}(X)\otimes\mathbb{C}^{n}$, which implies that the fibers $T(k)$ are independent of $k$. Henceforth, we will assume $T(k)\equiv T$. Although it is possible to generalize the analysis to a fully $k$-dependent $\Theta$, this extension lies beyond the scope of the present work. Straightforward computations then show that the commutation relation $TH=HT$ at the level of fibers reads
@@
Let us return to the fiber Hamiltonians. In most physically relevant models, as functions of the quasi-momentum $k$, these constitute an analytic family of operators in $L^{2}(\mathcal{W})\otimes\mathbb{C}^{n}$ in the sense of Kato [kato2013perturbation], and the spectrum of the full Hamiltonian satisfies $\sigma(H)=\bigcup_{k\in\mathbb{R}^{d}}\sigma(H(k))$ [reed1978iv]. Typically, $\sigma(H)$ consists of disjoint closed intervals, referred to as Bloch bands, which can be used, as shown by the following proposition, to construct projection-valued maps, which will be the main object of study in the following sections.
@@
Let $H(k)$ be a family of self-adjoint operators on a Hilbert space $\mathcal{H}$, and let $\gamma$ be a smooth, closed complex curve encircling exactly one Bloch band $\Omega$. Then
@@
defines a family of projection-valued maps satisfying $P_{\Omega}(k)=\chi_{\Omega}(H(k))$, where $\chi_{\Omega}$ is the characteristic function of $\Omega$. Moreover, $P_{\Omega}(k)$ inherits the same regularity in $k$ as $(H(k)-z)^{-1}$.
@@
Observe that if $f(w)=\frac{i}{2\pi}\int_{\gamma}\frac{1}{w-z}dz$, then $P_{\Omega}(k)=f(H(k))$. The integral computes the winding number of $\gamma$ around $w$, which is 1 if $w$ lies inside $\gamma$ and 0 otherwise. Hence, for the interval $\Omega$ enclosed by $\gamma$, we have $f=\chi_{\Omega}$, and thus
@@
An immediate corollary is that, for a sufficiently regular hamiltonian $H(k)$, the rank of $P_{\Omega}(k)$ is constant, since $\operatorname{tr}(P_{\Omega}(k))=\dim\operatorname{Ran}(P_{\Omega}(k))\equiv m\in\mathbb{N}$ defines a continuous map into a discrete set. Moreover, this family of projectors is again $\tau$-covariant, and in the time reversal symmetric setting it inherits a symmetry constraint from Equation (3):
$P_{\Omega}(k+\lambda^{*})=\tau_{\lambda^{*}}P_{\Omega}(k)\tau_{\lambda^{*}}^{-1},\quad TP_{\Omega}(k)=P_{\Omega}(-k)T,\quad\forall k\in\mathbb{B},\ \lambda^{*}\in\Lambda^{*}.$
Although the $\tau$-covariance of this family of projections is essential for defining the topological invariants we aim to study, the associated pseudoperiodicity can be cumbersome to handle. To address this, we employ the following construction, first introduced in [cornean2016construction]. We begin by considering a lattice basis $\lambda_{1}^{*},\lambda_{2}^{*},\dots,\lambda_{d}^{*}$ for $\Lambda^{*}$ and the corresponding unitary operators $\tau_{\lambda_{j}^{*}}(\phi)(y)=e^{i\lambda_{j}^{*}\cdot y}\phi(y)$ acting on $\mathcal{H}$. Using the spectral theorem for unitary operators, we find $d$ self-adjoint operators $L_{1},\dots,L_{d}$ such that $\tau_{\lambda_{j}^{*}}=e^{iL_{j}}$. Since they commute with each other, they can be chosen to satisfy $L_{a},L_{b}]=0$ for all $a,b=1,\dots,d$.

### v2

@@ Abstract.
Our main contributions are threefold. First, we establish a splitting theorem, proving that any such projection-valued map admits a splitting into two projection-valued maps that are related to each other via time-reversal symmetry. Second, we provide a complete homotopy classification theorem for these maps, thereby clarifying their topological structure. Third, by means of the previous analysis, we connect the Fu–Kane–Mele index to the Chern number of one of the factors in the previously-mentioned decomposition, which in turn allows to exhibit how the $\mathbb{Z}_{2}$-valued topological obstruction to constructing a periodic and smooth Bloch frame for the projection-valued map, measured by the Fu–Kane–Mele index, can be concentrated in a single pseudo-periodic Kramers pair.
@@ Introduction
Class AII corresponds to systems endowed with an antiunitary time-reversal symmetry $\Theta$ satisfying $\Theta^{2}=-1$. For crystalline and insulating systems, described by a periodic Hamiltonian with a spectral gap, one associates the so-called Fermi projection $P$, which defines the occupied-energies subspace. As reviewed below, this in turn gives rise to the Bloch bundle $E\to\mathbb{T}^{d}$ over the Brillouin torus $\mathbb{T}^{d}$, carrying an induced antiunitary equivariant structure. Fiber bundles subject to this type of symmetry are called “Quaternionic”-vector bundles, and have became object of independent geometric interest, beyond the specific case considered in this paper, due to their peculiar topological structure (see, e.g., [de2015classification, bunk2019topologicalinsulatorskanemeleinvariant]).
@@
As shown in [panati2007triviality, monaco2015symmetry], in two-dimensional systems (to which we restrict for the rest of this Introduction) the presence of $\Theta$ guarantees the existence of a global smooth trivialization of $E$, corresponding to the vanishing of the Chern number $\operatorname{Ch}(P)\in\mathbb{Z}$. However, such a trivialization need not respect the time-reversal symmetry itself, and the obstruction to the existence of a symmetric trivialization carries non-trivial topological content. This obstruction is quantified by a $\mathbb{Z}_{2}$-valued invariant, first introduced in [kane2005z]. While the original formulation of this invariant was already deeply topological in nature, several alternative formulations and interpretations have since been proposed, both in the physical and in the mathematical literature, see e.g. [graf2013bulk, de2015classification, fiorenza2016z, monaco2015symmetry, FuKane2006, MooreBalents2007]. For an overview of the relation among some of the different mathematical approaches, we refer the reader to [Cornean_2017].
@@
Following the approach initiated in [fiorenza2016z], the $\mathbb{Z}_{2}$ invariant can be interpreted as the obstruction to the existence of a globally symmetric trivialization of the Bloch bundle. In position space, the analog of a globally smooth trivialization of the Bloch bundle is the existence of exponentially localized Wannier bases. Such bases have a clear physical interpretation as representing localized electronic orbitals, and their existence is intimately related to the insulating nature of the system. The requirement that a Wannier basis is symmetric under time-reversal symmetry allows to implement the symmetry at the level of such orbitals, which can be useful when these are used, for example, to derive symmetric tight-binding models from continuum Hamiltonians [marzari2012maximally, marrazzo2024wannier]. In view of their relevance to computational solid state physics, we comment in Section 5.3 on how the topological obstruction to localization of time-reversal symmetric Wannier orbitals can be avoided if one relaxes the notion of orthonormality to that of forming a Parseval frame (compare [cornean2019parseval]).
@@
If, in addition to $\Theta$, the system possesses a unitary involutive symmetry $S$ that is compatible with $\Theta$ —corresponding, for example, to conservation of one component of spin— then the topological invariant persists and admits a simple physical interpretation as the spin Chern number [Sheng_2006, Prodan_2009]
$\operatorname{Ch}_{S}(P):=\frac{\operatorname{Ch}\big(P\tfrac{S+\mathds{1}}{2}\big)-\operatorname{Ch}\big(P\tfrac{S-\mathds{1}}{2}\big)}{2}.$
As shown by Prodan [Prodan_2009], the spin Chern number is stable under small perturbations, remaining well defined even when the symmetry $S$ is slightly broken – precisely, as long as the operator $PSP$ has a gap around 0. Moreover, the same work discusses how, upon changing the spin-type symmetry $S$, the value of the spin Chern number can change only modulo $2$, making it a natural candidate for the invariant associated with the AII symmetry class.
The role of the symmetry $S$ in computing the Spin-Chern number is essentially to split the Fermi projection into two components exchanged by $\Theta$:
(1) $P=P^{+}+P^{-}\,,\qquad P^{\pm}:=\frac{1}{2}\big(P\pm PS\big),\qquad\Theta P^{+}\Theta^{-1}=P^{-}.$
The spin Chern number then coincides with the Chern number of either component (which necessarily agree with each other). We will refer to such a decomposition of $P$ as a symmetric splitting. The existence of a symmetric splitting may be viewed as a weaker form of the existence of a spin–type symmetry. In the presence of such a splitting, one can still define an integer-valued invariant, which moreover enjoys the same robustness and locality properties as the Chern number. This perspective has already been partially explored in the physics literature, for instance in [Gilardoni_2022, Ba__2024].
In the present work, we prove that, for periodic systems endowed with fermionic time-reversal symmetry it is always possible to find a symmetric splitting. We show that the topological characterization of such systems is captured by the Chern number, modulo $2$, of one component of the splitting, or equivalenly by the Spin-Chern number of $P$ with respect to a spin-type operator taylored around the splitting. The $\mathbb{Z}_{2}$ nature of the invariant arises from the non-uniqueness of this splitting or of this spin-type symmetry.
Our proof of this result passes through the construction of an appropriate trivialization of the Bloch bundle, as mentioned above, which is compatible with time-reversal symmetry but possibly not with the full lattice periodicity. This construction unlocks also the possibility to rephrase the $\mathbb{Z}_{2}$ index as an invariant of the homotopy and unitary equivalence classes (and even of the Murray–von Neumann equivalence class; compare Definition 1.8) of the projection $P$, showing how the invariant is stable with respect to continuous deformations of the physical model – that is, of its Hamiltonian – which keep the spectral gap open, as well as with respect to unitary changes of representation which cannot affect the description of the physical properties of the system.
@@
These results provides a new perspective on the geometric and physical meaning of the Fu–Kane–Mele index: in particular, unlike other topological formulations, it admits a clear position-space counterpart, namely the Chern marker of one of the splitting projections. This position-space counterpart of the invariant paves the way for extending several results previously obtained in the spin-conserving case to the more general symmetry class AII, such as the addition of disorder and impurities in the crystal [bellissard1994noncommutative].
The paper is structured as follows. The first section is devoted time-reversal symmetric and periodic Hamiltonian, and how their gapped spectral eigenprojections allow to define time-reversal symmetric periodic projection-valued maps on the (Brillouin) torus. The second section revisits parallel transport in the context of continuous, rather than differentiable, projection-valued maps. The third section reviews some known results on projection-valued maps in absence of time-reversal symmetry, including the construction of continuous frames (i.e. trivializations) for them and how, in $d=2$, the Chern numbers obstructs the possibility to impose periodicity on the frame. The fourth section introduces the $\mathbb{Z}_{2}$-valued topological invariant for two-dimensional time-reversal symmetric projection-valued maps, which has a similar interpretation as a topological obstruction. The main results of the paper, regarding this $\mathbb{Z}_{2}$ invariant, are collected in the fifth section, and present its relation to symmetric splittings, the Spin-Chern number, and the completeness of the invariant as a marker for the different equivalences discussed above. Finally, the appendix presents a series of classical results in basic algebraic topology that are used throughout the paper.
Acknowledgements.
A. F. and G. P. would like to thank the organizers of the Quantum Mathematics @ PoliMI Intensive Period for the kind hospitality. The Authors also thank Marco Valerio D’Agostino and Gianluca Panati for valuable discussions on topics related to the paper.
D. M. and G. P. gratefully acknowledge financial support from Sapienza Università di Roma within Progetto di Ricerca di Ateneo 2023 and 2024, as well as from MUR–Italian Ministry of University and Research and Next Generation EU within PRIN 2022AKRC5P “Interacting Quantum Systems: Topological Phenomena and Effective Theories”. The Authors gratefully acknowledge financial support from PNRR MUR under Project No. PE0000023-NQSTI.
This work has been carried out under the auspices of the GNFM-INdAM (Gruppo Nazionale per la Fisica Matematica — Istituto Nazionale di Alta Matematica).
1. From (symmetric) periodic Hamiltonians to (symmetric) projection-valued maps
@@
In this section, we outline the standard procedure that allows one to associate a projection-valued map to a periodic Hamiltonian, following the construction reviewed in [panati2007triviality, panati2013bloch] to which we refer the reader for further details. We consider a quantum particle moving in a condensed matter system, whose configuration space $X\subset\mathbb{R}^{d}$ exhibits a periodic structure with respect to a lattice $\Lambda\simeq\mathbb{Z}^{d}$, commonly referred to as the Bravais lattice. Explicitly, this periodicity implies $X=\lambda+X$ for every $\lambda\in\Lambda$. Two types of models are typically used in this context: continuum models, in which $X=\mathbb{R}^{d}$, and tight-binding models, where $X$ is replaced by the discrete set of atomic positions and the quotient $X/\Lambda$ is finite. When $n$ additional internal degrees of freedom, such as spin, are present, the configuration space is naturally extended to $X\times\{1,\ldots,N\}$, and the corresponding Hamiltonian $H$ describing the system acts on $L^{2}(X)\otimes\mathbb{C}^{N}$. Furthermore, the system is assumed to be endowed with a fermionic time-reversal symmetry, i.e., an anti-unitary operator $\Theta$ acting on $L^{2}(X)\otimes\mathbb{C}^{N}$ such that $\Theta^{2}=-\mathds{1}$, commuting with the Hamiltonian $H$. The discussion therefore puts on the same footing both continuum models ($X=\mathbb{R}^{d}$), described by Hamiltonians with spin-orbit coupling of Rashba type [rashba2005spin], and discrete, tight-binding models, like the paradigmatic Kane–Mele Hamiltonian [Kane_2005, kane2005z]. For the sake of presentation, we’ll start from setting up the mathematical framework in continuum models: compare Remark rmk:discrete for the minor changes that take place in the discrete setting.
@@
where $(T_{\lambda}\phi)(x)=\phi(x-\lambda)\in\mathbb{C}^{N},$ for all $\lambda\in\Lambda$. Correspondingly, the fermionic time-reversal symmetry operator $\Theta$ is required to be itself translation-invariant, i.e., to commute with the lattice translations $T_{\lambda}$.
A powerful tool to study the spectral properties of such periodic Hamiltonians is the Bloch-Floquet-Zak transform [kuchment2016overview, panati2013bloch], which allows one to decompose the Hilbert space into fibers labeled by the quasi-momentum $k$. For any sufficiently rapidly decaying $\phi\in L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{N}$, the transform is defined as the following function depending on $k\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$:
@@
$(\mathcal{U}_{\mathrm{BFZ}}\phi)_{k}(y)\in\left\{f\in L^{2}_{\rm loc}(\mathbb{R}^{d},dk,\mathcal{H}_{f})\ \middle|\ f(k+\lambda^{*})=\tau_{\lambda^{*}}f(k)\right\}=:\mathcal{H}_{\tau}\simeq\int_{\mathbb{B}}^{\oplus}dk\,\mathcal{H}_{f},$
where $\mathcal{H}_{f}$ is the space of $\mathbb{C}^{N}$-valued, $\Lambda$-periodic, locally square-integrable functions on $\mathbb{R}^{d}$, and $\tau_{\lambda^{*}}:\mathcal{H}_{f}\to\mathcal{H}_{f}$ acts as $\tau_{\lambda^{*}}(g)(y)=e^{i\lambda^{*}\cdot y}g(y)$. By also choosing a periodicity cell for $\Lambda$, the Wigner-Seitz cell $\mathbb{W}$, one can further identify $\mathcal{H}_{f}$ with $L^{2}_{\rm per}(\mathbb{W})\otimes\mathbb{C}^{N}$, the space of $\mathbb{C}^{N}$-valued square-integrable functions on $\mathbb{W}$ with $\Lambda$-periodic boundary conditions. A classical result, which can be traced back to [bloch1929quantenmechanik, zak1964magnetic] (compare [reed1978iv]), reads:
@@
The operator $\mathcal{U}_{\mathrm{BFZ}}$ defined above extends uniquely to a unitary operator
$\mathcal{U}_{\mathrm{BFZ}}:L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{N}\longrightarrow\int_{\mathbb{B}}^{\oplus}dk\,L^{2}_{\rm per}(\mathbb{W})\otimes\mathbb{C}^{N}.$
When $H$ acting in $L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{N}$ commutes with the translation operators $T_{\lambda}$, $\lambda\in\Lambda$, the transformed operator $\mathcal{U}_{\mathrm{BFZ}}\,H\,\mathcal{U}_{\mathrm{BFZ}}^{-1}$ is decomposable: there exists a family of self-adjoint fiber operators $\{H(k)\}_{k\in\mathbb{R}^{d}}$ such that
$\big[\mathcal{U}_{\mathrm{BFZ}}\,H\,\mathcal{U}_{\mathrm{BFZ}}^{-1}f\big]_{k}=H(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{R}^{d},$
@@
The same procedure can be applied to $\Theta$, since it commutes with the translation operators. Consequently, there exists a family of anti-unitary operators $T(k)$ acting on $\mathcal{H}_{\mathrm{f}}$ such that
$\left[\mathcal{U}_{\mathrm{BFZ}}\,\Theta\,\mathcal{U}_{\mathrm{BFZ}}^{-1}(f)\right]_{k}=T(k)f_{-k},\quad\forall f\in\mathcal{H}_{\tau},\ k\in\mathbb{R}^{d}.$
@@
$T(k+\lambda^{*})=\tau_{\lambda^{*}}T(k)\tau_{\lambda^{*}},\quad\forall k\in\mathbb{R}^{d},\ \lambda^{*}\in\Lambda^{*}.$
In most physical models, $\Theta$ acts only on the second factor of the tensor product $L^{2}(X)\otimes\mathbb{C}^{N}$, which implies that the fibers $T(k)$ are independent of $k$. Henceforth, we will assume $T(k)\equiv T$. Although it is possible to generalize the analysis to a general $\Theta$, this extension lies beyond the scope of the present work. Straightforward computations then show that the commutation relation $TH=HT$ at the level of fibers reads
@@
Let us return to the fiber Hamiltonians. In most physically relevant models, the mapping $k\mapsto H(k)$ constitutes an analytic family of operators in $\mathcal{H}_{\mathrm{f}}$ in the sense of Kato [kato2013perturbation], meaning in particular that the resolvent operators $(H(k)-z)^{-1}$ are bounded operators depend analytically on the momentum $k$; moreover, the spectrum of the full Hamiltonian satisfies $\sigma(H)=\bigcup_{k\in\mathbb{R}^{d}}\sigma(H(k))$ [reed1978iv]. Typically, $\sigma(H)$ consists of disjoint closed intervals, referred to as Bloch bands, which can be used, as shown by the following proposition, to construct projection-valued maps, which will be the main object of study in the following sections.
@@
Let $H(k)$ be a family of fiber Hamiltonians, and let $\Omega$ be a set of Bloch bands that is spectrally separated from the rest of the spectrum. Let $\gamma$ be a smooth, closed, positively oriented contour contained in the resolvent set of $H(k)$ for all $k$, enclosing $\Omega$ and no other part of the spectrum. Then
@@
defines a family of projections on $\mathcal{H}_{\mathrm{f}}$ satisfying $P_{\Omega}(k)=\chi_{\Omega}(H(k))$, where $\chi_{\Omega}$ denotes the spectral projection onto $\Omega$. Moreover, the map $k\mapsto P_{\Omega}(k)$ inherits regularity in $k$, $\tau$-covariance, and (if present) time–reversal symmetry from the map $k\mapsto(H(k)-z)^{-1}$.
@@
Observe that if $f(w)=\frac{i}{2\pi}\int_{\gamma}\frac{1}{w-z}dz$, then $P_{\Omega}(k)=f(H(k))$. The integral computes the winding number of $\gamma$ around $w$, which is 1 if $w$ lies inside $\gamma$ and 0 otherwise. By assumption, the spectrum of $H(k)$ inside $\gamma$ is contained in $\Omega$. Therefore,
@@
An immediate corollary is that the rank of $P_{\Omega}(k)$ is constant —say, equal to $m\in\mathbb{N}$—, since $k\mapsto\operatorname{tr}(P_{\Omega}(k))=\dim\operatorname{Ran}(P_{\Omega}(k))$ defines a continuous map into a discrete set. Henceforth, we denote by $\operatorname{Proj}_{m}(\mathcal{H}_{\mathrm{f}})$ the space of rank-$m$ on the fiber Hilbert space $\mathcal{H}_{\mathrm{f}}$.
Let us spell out $\tau$-covariance and the time-reversal symmetry constraint inherited from Equations (2) and (3):
$P_{\Omega}(k+\lambda^{*})=\tau_{\lambda^{*}}P_{\Omega}(k)\tau_{\lambda^{*}}^{-1},\quad TP_{\Omega}(k)=P_{\Omega}(-k)T,\quad\forall k\in\mathbb{R}^{d},\ \lambda^{*}\in\Lambda^{*}.$
Although the $\tau$-covariance of this family of projections is essential for defining the topological invariants we aim to study, the associated pseudo-periodicity can be cumbersome to handle. To address this, we employ the following construction, first introduced in [cornean2016construction]. We begin by considering a lattice basis $\{\lambda_{1}^{*},\lambda_{2}^{*},\dots,\lambda_{d}^{*}\}$ for $\Lambda^{*}$ and the corresponding unitary operators $\tau_{\lambda_{j}^{*}}(\phi)(y)=e^{i\lambda_{j}^{*}\cdot y}\phi(y)$ acting on $\mathcal{H}$. Using the spectral theorem for unitary operators, we find $d$ self-adjoint operators $L_{1},\dots,L_{d}$ such that $\tau_{\lambda_{j}^{*}}=e^{iL_{j}}$. Since the $\tau_{\lambda_{j}^{*}}$’s commute with each other, they can be chosen to satisfy $L_{a},L_{b}]=0$ for all $a,b\in\{1,\dots,d\}$.